madalainerose:

In ΔXYZ, m∠X = 90° and m∠Y = 30°. In ΔTUV, m∠U = 30° and m∠V = 60°. Which is true about the two triangles? ΔXYZ ≅ ΔTUV ΔXYZ ≅ ΔVUT No congruency statement can be made because only two angles in each triangle are known. No congruency statement can be made because the side lengths are unknown.

Tranquility:

For two triangles to be congruent, both the angles and the sides have to be the same. For triangle XYZ, what is the missing angle if \(\angle X = 90^o\) and \(\angle Y = 30^o\) For triangle TUV, what is the missing angle if \(\angle U= 30^o\) and \(\angle V = 60^o\)

madalainerose:

i dont really understand but thank u @Tranquility

madalainerose:

but i also havent been in a math class in forever. ive been taught through Edgenuity.

Tranquility:

 Does the 30-60-90 triangle seem familiar?

madalainerose:

yes

Tranquility:

What would the missing angle be?

Tranquility:

For both the triangles

madalainerose:

lol um

madalainerose:

is it no congruency statement can be made because the side lengths are unknown?

Tranquility:

Do not spam the post @MioneCamp Last warning

Tranquility:

That's correct! @madalainerose Although the angles are the same for the triangle, we can't say it's congruent because we don't know anything about the side lengths

madalainerose:

yayyyyyy okay thank u @Tranquility

Tranquility:

Of course, no problem!